OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle with legs of lengths 9 and 12. If necessary, round your answer to two decimal places.

OpenStudy (anonymous):

use the pythagorean theorem a^2 + b^2 = c^2 so then a = 9 and b = 12 and solve for c

OpenStudy (anonymous):

there is one right angled triangle 3,4 and 5 - that slould give u a clue

OpenStudy (anonymous):

oh yeah.. okay

OpenStudy (anonymous):

the pythagorean theorem says that (leg)^2 x (leg)^2 = (hypotenuse)^2 given that the triangle is a right triangle

OpenStudy (anonymous):

dont you do the square root after?

OpenStudy (anonymous):

yes to find c since c is squared in the equation

OpenStudy (anonymous):

i got 15

OpenStudy (anonymous):

you can also find the equation using trig to find the angle the triangle makes and using that to find the hypothenuse

OpenStudy (anonymous):

15 is correct